OpenStudy (anonymous):

A bicyclist is riding on a path modeled by the function f(x) = 0.05(10x – x2), where x and f(x) are measured in miles. Find the rate of change of elevation when x = 3. (Round your answer to two decimal places.)

OpenStudy (anonymous):

Plug whatever x equals into the equation. \[f(x)=.05(10(3)-(3)^2)\] (assuming "x2" is "x^2")

OpenStudy (anonymous):

The derivative of (5/100)*(10x – x2) is:\[\frac{1}{20} (10-2 x) \]evaluated at x = 3 is:\[\frac{1}{5} \]Refer to the attachment